Circular No. 6319

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SCORPIUS X-1 M. van der Klis, Astronomical Institute ‘Anton Pannekoek’, University

of Amsterdam; J. Swank, W. Zhang, and K. Jahoda, Goddard Space Flight Center, NASA; E. Morgan and W. Lewin, Massachusetts Institute of Technology; B. Vaughan, California Institute of Technology; and J. van Paradijs, University of Alabama at Huntsville and University of Amsterdam, report: “We have observed quasiperiodic oscillations (QPOs) with frequencies between 1060 and 1130 Hz in two different 2- to 60-keV observations of Sco X-1 with the X-ray Timing Explorer satellite. In the first observation, on Feb. 14.470–14.492 UT, Sco X-1 was mostly in the lower normal branch of the Z track in the x-ray color-color diagram, with clear 6- to 7-Hz, normalbranch QPOs. Power spectra of two simultaneous, statistically-independent time series (resolution 16 and 62.5 μs, respectively) show a narrow QPO peak at 1062 ± 7 and 1077 ± 8 Hz, respectively, with a FWHM of ∼ 60 Hz and a deadtime-corrected, fractional, rms amplitude, r, of 0.74 ± 0.10 percent. In the second observation, on Feb. 18.203–18.234, Sco X-1 changed state several times, moving along the Z track between the lower normal and flaring branches, showing 7- and 16-Hz QPOs, respectively. Using 250-μs data, a very narrow QPO peak is seen in the normal branch at 1089.1 ± 0.5 Hz with FWHM 15 ± 3 Hz and r = 0.72 ± 0.06 percent. In the flaring branch, the peak is at 1133 ± 7 Hz, FWHM is roughly 40 Hz, and r = 0.49 ± 0.09 percent. In another segment of the same observation, 16- and 62-μs data taken in the normal branch again show similar high-frequency, comparatively-narrow peaks. All peak significances are between 3.5σ and 4.5σ (single trial). In both observations, there is evidence that the QPOs increase in frequency along the normal branch and into the flaring branch (i.e., with mass-transfer rate), and that part of the observed peak widths is due to this motion. In a third observation, when Sco X-1 was in the flaring branch, no significant kHz peak was seen. This is the fastest QPO yet seen in an accreting compact object. The high-QPO frequency, and its increase with mass-transfer rate, suggest that we may be seeing the keplerian frequency at the inner edge of the disk near the magnetospheric boundary, or its beat frequency with a slower (∼ 100 Hz) pulsar. The fact that the high-frequency QPOs have a similar dependence of strength and frequency on mass-transfer rate as the well-known 6- to 20-Hz QPOs in Sco X-1 indicates that the two phenomena might be related.”

1996 February 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green